If the length of a rectangle is 3 more than its width and the area of the rectangle is 154 square inches, then what are the dimensions of the rectangle?
Answer by math_helper(2461)


L = W + 3
LW = 154 sq in

Substitute for L in the LW=154:
(W+3)(W) = 154
W^2 + 3W - 154 = 0
(W + 14)(W - 11) = 0
W = 11 and/or W = -14 (discard W=-14 as W > 0 for a rectangle)
W=11 —> L=14

Ans: The rectangle is 11in x 14in

Sanity check: 11x14 = 154